40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Nov 8 03:36:44 2013, tara, DailyProgress, optic, redo- PMC path 
    Reply  Mon Nov 11 19:37:13 2013, tara, DailyProgress, optic, redo- PMC path 
Message ID: 1386     Entry time: Mon Nov 11 19:37:13 2013     In reply to: 1385
Author: tara 
Type: DailyProgress 
Category: optic 
Subject: redo- PMC path 

PMC path is back, I aligned the polarization of the input beam to the BB EOM for TTFSS. The visibility of PMC is now ~ 80%.

ELOG V3.1.3-